# Unsupervised Representation Learning from Sparse Transformation Analysis

### v1

@@ Appendix A Limitations and Future Work
One of the main limitations of this work in relation to prior work on independent component analysis and sparse coding is the unanswered question of ‘identifiability’ of latent factors in our model. Specifically, much of the early work on ICA was focused around answering to what extent the ground truth latent variables can be recovered after an unknown linear transformation [3, 93]. More recently, a number of papers have demonstrated the necessary conditions for these identifiability proofs to extend to non-linear transformations of the latent variables [94]. Interestingly, these necessary conditions often revolve around temporal structure such as time-step conditioning [95], or sparsity of transformations over time [5]. Our model, Sparse Transformation Analysis, clearly takes inspiration from these proofs by integrating these factors into a highly flexible generative modeling framework. Empirically, we demonstrate that STA does separate the ground-truth factors of variation from sequence data entirely unsupervised for a variety of datasets, including ‘real world’ datasets with more natural image statistics. We note that even in the case of models which have proved this identifiability, this is often only under strict assumptions about the true data distribution, and these assumptions are frequently seen to be invalidated on real data. These papers therefore similarly resort to measuring if their models are able to separate the underlying factors of variations in an empirical manner, as we have done in this paper. Furthermore, in our empirical analysis, we see that due to the flexibility endowed by the learned latent flows of our model, it is able to outperform models which indeed have these identifiability guarantees (such as [5]). That being said, in future work, we believe that it would undoubtedly be beneficial and of great interest to understand the exact necessary conditions for such a model to provably identify the generative factors. We believe such an understanding will likely be helpful for selecting optimal settings of the hyperparameters, such as the probability of switching on for the Bernoulli distribution or the scale parameter for the Laplace distribution, while similarly helping to guide the future development of unsupervised representation learning in general.
@@
Appendix B Proof of how the Hamilton-Jacobi equation solves Optimal Transport
Theorem B.1 (Benamou-Brenier Formula [90]).
@@ Appendix B Proof of how the Hamilton-Jacobi equation solves Optimal Transport
[…]
@@
$\begin{cases}\partial_{m}L=v-\nabla u=0\\ \partial_{u}L=\partial_{t}\rho+\nabla\cdot m=0\\ \partial_{\rho}L=-\frac{||m||^{2}}{2\rho^{2}}-\partial_{t}u=-\frac{1}{2}||v||^{2}-\partial_{t}u=0\\ \end{cases}$ (27)
where the first condition indicates that the velocity field $v(x,t)$ is given by the gradient $\nabla u(x,t)$, the second condition gives the continuity equation which holds in the sense of distributions, and the third condition yields the optimal solution for minimizing the Wasserstein distance — the Hamilton-Jacobi equation ($\partial_{t}u+\frac{1}{2}||\nabla u||^{2}=0$).
Appendix C Additional Implementation Details
C.1 Model Architectures
For the MLPs that parameterize the scalar potential $u({\bm{z}},t)$ and the divergence-free vector field ${\bm{r}}({\bm{z}})$, we use linear layers to encode the latent samples and use Tanh as the activation function. The sinusoidal positional embeddings [96] are used to embed the timestep $t$. For our variational auto-encoders, the encoder simply takes four stacked convolution layers with ReLU activation functions, while the decoder consists of four transposed convolution layers.
C.2 Data Sequences and Baselines
@@
For LatentFlow, PoFlow, and TVAE which are supervised baselines, we use their respective latent operators to move latent samples. For the unsupervised approaches, we carefully select the latent dimension that corresponds to the lowest equivariance error of a given transformation and perform a grid search to tune the traversal range in the interval $[-5,5]$. Since the vanilla VAE does not have any notion of learned latent transformations, when computing the equivariance error, we simply take it as a lower-bound baseline by setting the latent samples unchanged (i.e., ${\bm{z}}_{0}={\bm{z}}_{1}=\dots={\bm{z}}_{T}$).
C.3 Training Details
On MNIST [82], the training process lasts $50,000$ iterations. We only train the spike component in the first $20,000$ iterations and then integrate the slab component into the training for the rest of the iterations. On Shapes3D [91], we set the total training iterations to $100,000$ and split the iterations fifty-fifty into the two training stages. The batch size is set to $128$ and $64$ on MNIST and Shapes3D, respectively. We use the Adam optimizer with a learning rate of $41e{-}$ 4. The input images are of the size $28{\times}28$ on MNIST and of the size $64{\times}64$ on Shapes3D. The quantitative results are reported as mean $\pm$ standard deviation computer over $5$ runs with random initialization.
@@ C.3 Training Details
$\nabla u({\bm{z}},t)$ 303.81$\pm$5.13 271.74$\pm$4.56 286.52$\pm$4.96 -2118.49$\pm$3.09
$\nabla u({\bm{z}},t)+{\bm{r}}({\bm{z}})$ 281.32$\pm$4.71 230.93$\pm$5.02 292.85$\pm$4.58 -2107.65$\pm$2.27
Appendix D More Experimental Results
D.1 Disentanglement Metrics
As discussed in [28, 12], following $\beta$-VAE [18], there are many disentanglement metrics proposed for single-dimension traversal methods, such as DCI [97] and MIG [38]. These metrics assume that each latent dimension corresponds to one generative factor and manipulating these single dimensions would trigger distinct output transformations. The recent disentanglement methods [25, 28, 33] propose a more realistic disengagement setting: all the latent dimensions are perturbed by vector arithmetic for output variations. If one evaluates these vector-based disentanglement methods using the metrics designed for single-dimension manipulations, their scores would drop considerably and are not comparable. Nonetheless, certain disentanglement metrics such as the VP score [23] can be adopted for the evaluation of general disentanglement baselines as they do not pose any assumptions on the latent space. Instead, the VP metric takes a lightweight neural network to learn classifying a dataset of different image pairs $[{\bm{x}}_{0},{\bm{x}}_{T}]$ under the few-shot learning setting (i.e., only $1\%$ or $10\%$ of the dataset is used as the training set). The validation accuracy reflects the distinguishability of these learned traversal directions, which is a reasonable surrogate for the disentanglement score.
@@ D.1 Disentanglement Metrics
D.2 Impact of Divergence-free Vector Fields
@@
D.2.1 Traversal for More Steps
Fig. 15 displays the traversals of our learned flow fields for more time points (longer sequences). We see that our STA still allows for smooth interpolation across timesteps. Notice that the traversal step is even larger than the maximal sequence length of Falcol3D and Issac3D [14], but the latent flows keep smooth transitions during time evolution.

### v2

@@ Appendix A Limitations and Future Work
Limitations. One of the main limitations of this work in relation to prior work on independent component analysis and sparse coding is the unanswered question of ‘identifiability’ of latent factors in our model. Specifically, much of the early work on ICA was focused around answering to what extent the ground truth latent variables can be recovered after an unknown linear transformation [comon1994independent, 1306473]. More recently, a number of papers have demonstrated the necessary conditions for these identifiability proofs to extend to non-linear transformations of the latent variables [hyvarinen2023nonlinear]. Interestingly, these necessary conditions often revolve around temporal structure such as time-step conditioning [pmlr-v89-hyvarinen19a], or sparsity of transformations over time [klindt2021towards]. Our model, Sparse Transformation Analysis, clearly takes inspiration from these proofs by integrating these factors into a highly flexible generative modeling framework. Empirically, we demonstrate that STA does separate the ground-truth factors of variation from sequence data entirely unsupervised for a variety of datasets, including ‘real world’ datasets with more natural image statistics. We note that even in the case of models which have proved this identifiability, this is often only under strict assumptions about the true data distribution, and these assumptions are frequently seen to be invalidated on real data. These papers therefore similarly resort to measuring if their models are able to separate the underlying factors of variations in an empirical manner, as we have done in this paper. Furthermore, in our empirical analysis, we see that due to the flexibility endowed by the learned latent flows of our model, it is able to outperform models which indeed have these identifiability guarantees (such as [klindt2021towards]). That being said, in future work, we believe that it would undoubtedly be beneficial and of great interest to understand the exact necessary conditions for such a model to provably identify the generative factors. We believe such an understanding will likely be helpful for selecting optimal settings of the hyperparameters, such as the probability of switching on for the Bernoulli distribution or the scale parameter for the Laplace distribution, while similarly helping to guide the future development of unsupervised representation learning in general.
Generating High-Resolution Videos. When dealing with high-resolution videos, the complexity of detailed objects and diverse backgrounds substantially increases, making it extremely challenging to effectively capture intricate motions or transformations in the highly compressed latent space. To address these challenges, incorporating additional sophisticated feature extraction methods (such as key-point tracking) could be beneficial, as demonstrated in previous motion tracking or synthesis work [sun2021multi, tulyakov2018mocogan]. Thus, the primary obstacles lie in designing powerful motion feature extraction methods and identifying meaningful latent representations that can robustly encode complex motion dynamics at higher resolutions into lower dimensions.
Integration into Diffusion Models. Another critical obstacle is that standard diffusion models do not inherently define a semantically meaningful, highly compressed latent space. To integrate our methodology into diffusion-based models, we must first carefully identify an appropriate latent representation. One potential approach is leveraging the “h-space” proposed by Kwon et al. [kwon2022diffusion], which uses the bottleneck features from the score-prediction U-Net as a latent representation. However, empirical validation is still necessary to confirm whether such latent spaces are sufficiently powerful to capture transformations effectively compared to latent spaces explicitly constructed by GANs or VAEs.
Appendix B Identifiability of Sparse Transformation Analysis
We provide a formal argument supporting the identifiability of our sparse transformation model under mild assumptions. Specifically, we show that the latent vector fields and sparse transformation coefficients are identifiable (up to permutation and scaling) given observed sequential data.
B.1 Assumptions
Let ${\bm{x}}_{t}\in\mathbb{R}^{D}$ be the observation at time $t$, and ${\bm{z}}_{t}\in\mathbb{R}^{d}$ be the latent code. The latent evolution follows:
${\bm{z}}_{t}={\bm{z}}_{t-1}+\sum_{k=1}^{K}{\bm{g}}_{t}^{k}{\bm{v}}_{k}({\bm{z}}_{t-1}),$ (23)
where ${\bm{v}}_{k}(z)=\nabla{\bm{u}}_{k}(z)+{\bm{r}}_{k}(z)$ is the $k$-th vector field (Helmholtz decomposition), and ${\bm{g}}_{t}\in\mathbb{R}^{K}$ is a sparse coefficient vector. We make the following assumptions:
  • A1 (Smooth Decoding): The observation mapping $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ is a diffeomorphism. Thus, ${\bm{x}}_{t}=f({\bm{z}}_{t})$ and ${\bm{z}}_{t}$ can be recovered from ${\bm{x}}_{t}$.
  • A2 (Sparse Composition): The transformation coefficients ${\bm{g}}_{t}$ are $s$-sparse: ${\bm{g}}_{t}||_{0}\leq s$ with $s\ll K$.
  • A3 (Vector Field Independence): For almost every $\bm{z}$}, the set $\{{\bm{v}}_{k}({\bm{z}})\}_{k=1}^{K}$ is linearly independent.
  • A4 (Sufficient Support Diversity): The dataset contains a set of time steps $t$ such that the support patterns ${\mathbf{S}}_{t}=\text{supp}({\bm{g}}_{t})$ vary and sufficiently span all combinations of up to $s$ active components.
B.2 Theorem (Identifiability)
Under assumptions A1–A4, the set of vector fields $\{{\bm{v}}_{k}(\cdot)\}_{k=1}^{K}$ and the sparse transformation coefficients $\{{\bm{g}}_{t}\}_{t=1}^{T}$ are identifiable up to permutation and scaling.
B.3 Proof
From A1, we can invert the observations to obtain the latent sequence $\{{\bm{z}}_{t}\}$. Define the latent displacement:
$\delta{\bm{z}}_{t}:={\bm{z}}_{t}-{\bm{z}}_{t-1}=\sum_{k\in{\mathbf{S}}_{t}}{\bm{g}}_{t}^{k}{\bm{v}}_{k}({\bm{z}}_{t-1}),$ (24)
where ${\mathbf{S}}_{t}=\text{supp}({\bm{g}}_{t})$ and $|{\mathbf{S}}_{t}|\leq s$.
At each $t$, $\delta{\bm{z}}_{t}$ is a sparse linear combination of the set $\{{\bm{v}}_{k}({\bm{z}}_{t-1})\}$. Under A3 and A4, we obtain a sequence of sparse coding problems:
[…]
where ${\mathbf{V}}({\bm{z}}_{t-1})=[{\bm{v}}_{1}({\bm{z}}_{t-1}),\dots,{\bm{v}}_{K}({\bm{z}}_{t-1})]\in\mathbb{R}^{d\times K}$. From the theory of sparse dictionary learning [donoho2003optimally], if a dictionary $\mathbf{V}$} has columns that are linearly independent and the sparsity level $s$ satisfies $s<\text{spark}({\mathbf{V}})/2$, then for every vector $\delta{\bm{z}}$ that admits a $s$-sparse representation over $\mathbf{V}$}, this representation is unique. The spark of a matrix is the smallest number of columns that are linearly dependent. If $\mathbf{V}$} has full column rank and no subset of $2s$ or fewer columns is linearly dependent, then the sparse coefficients and the dictionary can be uniquely recovered (up to permutation and scaling) from sufficiently many samples.
Moreover, under the assumption A4, the dataset provides a diverse set of support patterns ${\mathbf{S}}_{t}$, ensuring that the different combinations of active vector fields are sufficiently sampled. This property is crucial for the joint recovery of the dictionary ${\mathbf{V}}({\bm{z}})$ and sparse codes ${\bm{g}}_{t}$ using methods akin to K-SVD [aharon2006k] or ER-SpUD [spielman2012exact]. When such diversity holds across $\bm{z}$}, the function-valued dictionary ${\bm{v}}_{k}({\bm{z}})$} can be recovered pointwise, as each ${\bm{z}}_{t-1}$ provides local linear constraints.
Therefore, the latent transformation structure – both vector fields $\{{\bm{v}}_{k}\}$ and sparse codes $\{{\bm{g}}_{t}\}$ – is identifiable up to permutation and scaling.
B.4 Identifiability of Helmholtz Components
This identifiability result also applies to the individual components of the Helmholtz decomposition. Specifically, the divergence-free components ${\bm{r}}_{k}$} (vorticity) and the curl-free components $\nabla{\bm{u}}_{k}$} (potential flows) are each identifiable up to permutation and scaling, provided that they appear independently or in sufficiently varied combinations within the support patterns. Since the decomposition is additive and the dictionary learning problem is posed over the combined flows ${\bm{v}}_{k}=\nabla{\bm{u}}_{k}+{\bm{r}}_{k}$, the linear independence and sparse excitation across time guarantee that both components can be separately identified as long as they do not systematically co-occur. In practice, this holds because different transformation types (e.g., rotation vs. scaling) are encoded by different structural priors and appear in distinct contexts.
@@ B.4 Identifiability of Helmholtz Components
Appendix C Proof of how the Hamilton-Jacobi equation solves Optimal Transport
Theorem C.1 (Benamou-Brenier Formula [benamou2000computational]).
@@ Appendix C Proof of how the Hamilton-Jacobi equation solves Optimal Transport
$\begin{gathered}W_{2}(\mu_{0},\mu_{1})^{2}=\min_{\rho,v}\Big\{\int\int\frac{1}{2}\rho(x,t)|v(x,t)|^{2}\\ \mathop{dx}\mathop{dt}:\frac{\mathop{d}\rho(x,t)}{\mathop{dt}}=-\nabla\cdot(v(x,t)\rho(x,t))\Big\}\end{gathered}$ (26)
@@
[…]
Appendix D Additional Implementation Details
D.1 Model Architectures
For the MLPs that parameterize the scalar potential $u({\bm{z}},t)$ and the divergence-free vector field ${\bm{r}}({\bm{z}})$, we use linear layers to encode the latent samples and use Tanh as the activation function. The sinusoidal positional embeddings [vaswani2017attention] are used to embed the timestep $t$. For our variational auto-encoders, the encoder simply takes four stacked convolution layers with ReLU activation functions, while the decoder consists of four transposed convolution layers.
D.2 Data Sequences and Baselines
@@
For LatentFlow, PoFlow, and TVAE which are supervised baselines, we use their respective latent operators to move latent samples. For the unsupervised approaches, we carefully select the latent dimension that corresponds to the lowest equivariance error of a given transformation and perform a grid search to tune the traversal range in the interval [5,5][-5,5]. Since the vanilla VAE does not have any notion of learned latent transformations, when computing the equivariance error, we simply take it as a lower-bound baseline by setting the latent samples unchanged (i.e., ${\bm{z}}_{0}={\bm{z}}_{1}=\dots={\bm{z}}_{T}$).
D.3 Training Details
On MNIST [lecun1998mnist], the training process lasts $50,000$ iterations. We only train the spike component in the first $20,000$ iterations and then integrate the slab component into the training for the rest of the iterations. On Shapes3D [3dshapes18], we set the total training iterations to $100,000$ and split the iterations fifty-fifty into the two training stages. The batch size is set to $128$ and $64$ on MNIST and Shapes3D, respectively. We use the Adam optimizer with a learning rate of 1e41e{-}4. The input images are of the size $28{\times}28$ on MNIST and of the size $64{\times}64$ on Shapes3D. The quantitative results are reported as mean $\pm$ standard deviation computer over $5$ runs with random initialization.
@@ D.3 Training Details
[…]
Appendix E More Experimental Results
E.1 Disentanglement Metrics
As discussed in [Tzelepis_2021_ICCV, song2023flow], following $\beta$-VAE [higgins2016beta], there are many disentanglement metrics proposed for single-dimension traversal methods, such as DCI [eastwood2018framework] and MIG [chen2018isolating]. These metrics assume that each latent dimension corresponds to one generative factor and manipulating these single dimensions would trigger distinct output transformations. The recent disentanglement methods [shen2021closed, Tzelepis_2021_ICCV, song2023householder] propose a more realistic disengagement setting: all the latent dimensions are perturbed by vector arithmetic for output variations. If one evaluates these vector-based disentanglement methods using the metrics designed for single-dimension manipulations, their scores would drop considerably and are not comparable. Nonetheless, certain disentanglement metrics such as the VP score [zhu2020learning] can be adopted for the evaluation of general disentanglement baselines as they do not pose any assumptions on the latent space. Instead, the VP metric takes a lightweight neural network to learn classifying a dataset of different image pairs ${\bm{x}}_{0},{\bm{x}}_{T}]$ under the few-shot learning setting (i.e., only $1\%$ or $10\%$ of the dataset is used as the training set). The validation accuracy reflects the distinguishability of these learned traversal directions, which is a reasonable surrogate for the disentanglement score.
TABLE X: VP Scores ($\%$) on MNIST.
@@ E.1 Disentanglement Metrics
E.2 Impact of Divergence-free Vector Fields
@@
TABLE XII: Approximation error of each divergence-free vector field on MNIST [lecun1998mnist].
$\nabla\cdot{\bm{r}}^{0}({\bm{z}})||_{2}^{2}$ $\nabla\cdot{\bm{r}}^{1}({\bm{z}})||_{2}^{2}$ $\nabla\cdot{\bm{r}}^{2}({\bm{z}})||_{2}^{2}$
0.017 0.032 0.028
Since our divergence-free vector fields $\nabla\cdot{\bm{r}}(\cdot)$ are enforced via a PINN loss, evaluating their approximation error is indeed important to verify their divergence-free property. Table XII reports the approximation errors for each divergence-free vector field trained on the MNIST dataset. The results indicate that our PINN effectively minimizes the divergence, achieving consistently small errors, thus validating that the learned vector fields closely approximate true divergence-free (rotational) fields.
@@
E.3 Traversal for More Steps
Fig. 15 displays the traversals of our learned flow fields for more time points (longer sequences). We see that our STA still allows for smooth interpolation across timesteps. Notice that the traversal step is even larger than the maximal sequence length of Falcol3D and Issac3D [nie2020semi], but the latent flows keep smooth transitions during time evolution.